OpenStudy (anonymous):

Factor completely: 3x2 − x − 4 (3x − 1)(x + 4) (3x + 4)(x − 1) (3x − 2)(x + 2) (3x − 4)(x + 1)

OpenStudy (anonymous):

the problem is =3x^2-x-4 =3x^2-(4-3)x-4 =3x^2-4x+3x-4 =x(3x-4)+1(3x-4) =(3x-4)(x+1)